Mrs. Evans is sixteen times as old as her daughter is, but in eight years she will only be four times as old as her daughter will be. How old is each now?
x=16y

Answer by macston(5194)

x=16y
M=mother; D=daughter
M=16D
M+8=4(D+8) Substitute for M.
16D+8=4D+32 Subtract 4D+8 from each side.
12D=24 Divide each side by 12.
D=2 ANSWER 1: The daughter is 2 years old.
M=16D=16(2)=32 ANSWER 2: The mother is 32 years old.
CHECK:
M+8=4(D+8)
32+8=4(2+8)
40=4(10)
40=40